# Exploiting the Layered Intrinsic Dimensionality of Deep Models for Practical Adversarial Training

### v1

@@ 3 Notation
Consider a deep neural network classifier $f_{\theta}$ with $n$ layers and parameters $\theta\!=\!\left\{\theta^{(i)}\right\}_{i=1}^{n}$. We denote by $f_{\theta^{[i,j]}}$ the nested transformation spanning layers $l_{i}$ to $l_{j}$, i.e., $f_{\theta^{[i,j]}}\!=\!f_{\theta^{(i)}}\!\circ\!\cdots\!\circ\!f_{\theta^{(j)}}$, and by $\bar{f}_{\theta^{[i,j]}}$ the transformation $f_{\theta^{[i,j]}}$ normalized to have zero mean and unit variance. Also, we denote by $\mathcal{\bm{X}}\in\mathbb{R}^{d\times|\mathcal{D}|}$ and $f_{\theta^{[1,l]}}(\mathcal{\bm{X}})\in\mathbb{R}^{d_{l}\times|\mathcal{D}|}$ the matrices obtained from stacking the input samples $\{x_{i}\}_{i=1}^{|\mathcal{D}|}$ and their transformations $\left\{f_{\theta^{[1,l]}}(x_{i})\right\}_{i=1}^{|\mathcal{D}|}$ vertically, for a given dataset $\mathcal{D}\!=\!\{x_{i},y_{i}\}_{i=1}^{|\mathcal{D}|}$.
The data manifold is described by the distribution of samples $\{x_{i}\}_{i=1}^{|\mathcal{D}|}$, i.e., samples in the high-density regions are on-manifold and samples in the low-density regions are off-manifold. Similarly, the $l^{\text{th}}$-layer manifold is characterized by the density of the normalized transformed samples, i.e., $\{\bar{f}_{\theta^{[1,l]}}(x_{i})\}_{i=1}^{|\mathcal{D}|}$. While these densities are unknown in practice, for simplicity we approximate the data and $l^{\text{th}}$-layer manifolds via the eigenspaces of the covariance matrices $\mathcal{\bm{X}}\mathcal{\bm{X}}^{T}$ and $\bar{f}_{\theta^{[1,l]}}(\mathcal{\bm{X}})\bar{f}_{\theta^{[1,l]}}(\mathcal{\bm{X}})^{T}$, respectively. We denote by $(U_{l},\Sigma_{l})$ the eigenspace of the $l^{\text{th}}$ layer, where $U_{l}\!\in\!\mathbb{R}^{d_{l}\times d_{l}}$ and $\Sigma_{l}\!\in\!\mathbb{R}^{d_{l}\times d_{l}}$ are the matrices of eigen-vectors and eigen-values, respectively. Since the eigenvectors $U_{l}$ are orthogonal (covariance matrices are symmetric), after normalization, the projection error on $k\in[1,d_{l}]$ eigenvectors is:
$e_{l}^{k}(x)=\bar{f}_{\theta^{[0,l]}}(x)-U_{l}^{k}{U_{l}^{k}}^{T}\bar{f}_{\theta^{[1,l]}}(x).$ (2)
The eigen-based data manifold of the $l^{\text{th}}$ layer is denoted by $k_{l}$ and is defined as the number of eigenvectors with the largest eigenvalues necessary for a reconstruction error $\gamma$ of the original inputs, i.e., $k_{l}\!=\!\min_{k}\left\{k\in[1,d_{l}]\Big{|}\sum_{x\in\mathcal{D}}\left\|e_{l}^{k}(x)\right\|_{2}\leq\gamma\right\}.$
An AE is $(\gamma,l)$-off-manifold, i.e., $(\gamma,l)$-OFM-AEs, if the projection error of its image at the input of the $l^{\text{th}}$ layer is larger than $\gamma$. Similarly, we define AEs as on-manifold (OMF-AEs) if $e_{l}^{k}\leq\gamma$.
The Intrinsic Dimension (ID) of the layer, estimated using the TwoNN method, is denoted as $I_{l}$. Furthermore, we denote the normalized ID of the layer by $I_{l}^{d}$ which is computed by dividing $I_{l}$ by $d_{l}$.
(a) BERT
(b) RoBERTa
(c) Vision Models
(d) dec-LLMs
Figure 3: The correlation between OFM/ONM AEs ratio (row 1) and the normalized ID (row 2). We use the average projection error ($e_{l}^{k}$) of AEs as a proxy for estimating the OFM/ONM AEs ratio and twoNN for ID estimation across layers of models including (a) BERT and (b) RoBERTa on AGNEWS, IMDB, and YELP datasets, (c) vision models on Cifar-10 dataset, and (d) dec-LLMs on HelpSteer dataset. Results show that enc-LLMs (BERT, RoBERTa) have decreasing ID trends unlike vision models and dec-LLMs that have increasing ID trends. These ID trends are reflected on reconstruction errors which show enc-LLMs have higher OFM/ONM AEs ratio at higher layers while vision models and dec-LLMs have at the initial layer.
4 Approach
4.1 Motivation and Overview
Our study is motivated by understanding the inherent difference in the generalization and robustness trends between adversarially trained vision and enc-LLMs. Specifically, (1) robustness improvement is consistently more significant in vision models [38, 4], and (2) generalization deteriorates in vision models, while it remains unchanged or even improves in enc-LLMs [61, 80].
We hypothesize that the underlying phenomenon is related to the distribution of OFM-AEs across the layers. Specifically, we speculate that in the first layer where conventional AT is applied, AEs tend to be more on-manifold for enc-LLMs, whereas they are more off-manifold for vision models and dec-LLMs. This would explain the results on generalization and robustness trends since it is established that higher proportions of on/off-manifold AEs lead, respectively, to better generalization/robustness [35, 55, 2, 31].
We study the relationship between the ID and the proportion of OFM/ONM AEs across different layers of various deepnet models, including two enc-LLMs (BERT and RoBERTa), three vision models (ViT [17], DeiT [58], VGG-13 [52]), and three dec-LLMs (LLaMA-2-7b [59], Vicuna [77], Mistral-7b [24]) on several dataset (refer to Appendix B for details). Our observation in the first row of Figure 3 shows that, for two enc-LLMs (BERT and RoBERTa), the proportion of OFM-AEs increases with the layer depth, in contrast to the trend observed in vision models and dec-LLMs. The empirical results confirm that normalized ID, $I_{l}^{d}$, follows the OFM/ONM ratio of AES since if the embedding space of the layer, $d_{l}$, is much greater than the ID of the layer, $I_{l}$, there is a higher probability for the samples to be OFM-AEs. In line with this, while $I_{l}^{d}$ decreases for the enc-LLMs, it increases for vision models [^2] Our findings are consistent with earlier studies that reveal a ”hunchback shape” —initially increasing and then decreasing– in the ID behavior of vision models [5]. As convolutional neural networks decrease in dimensionality $d_{l}$, the normalized ID, $I_{l}^{d}$, follows an increasing trend. and dec-LLMs. We also observe that estimated IDs and measured projection errors show a strong correlation across all models, with very few exceptions. (Refer to Appendix B for a further discussion on this issue.)
Inspired by these findings, we propose an approach for improving robustness over standard AT, via perturbing an intermediate layer $l^{*}$ with a low IDs. We show that our AT approach, Scalable Manifold Aware Adversarial Training (Fig. 2), significantly improves scalability by a factor $\mathcal{O}\Big{(}P(n-l^{*})\Big{(}\max\!\Big{(}\{d_{l_{i}}|l_{i}\in[1,l^{*}-1]\}\Big{)}-d_{l^{*}}\Big{)}\Big{)}$ for an $n$-layered deepnet under a $P$-step PGD attack.
4.2 SMAAT: Scalable Manifold Aware Adversarial Training
We propose SMAAT an adversarial training algorithm that aims at improving the robustness and scalability of AT by generating AEs in the layer leading to the highest proportion of OFM instead of from the input layer as classically. Specifically, given a pretrained model $f_{\theta}$, SMAAT intentionally generates a higher proportion of OFM-AEs to enhance robustness. Formally, SMAAT solves the augmented AT objective:
@@ 4.2 SMAAT: Scalable Manifold Aware Adversarial Training
While the manifold conjecture refers to input space AEs, we relax it to encompass intermediate layers as well, i.e., we chose the perturbation $\delta$ that results in an on-/off-manifold AE at any layer across the deepnet. Intuitively, AEs that are off or on the transformed data manifold in any layer also affect robustness. The relaxed objective becomes:
[…]
Note that solving the above objective using the method of Lagrangian Multipliers is possible, but it would require an approximation of the manifold to characterize OFM-AEs. Such approximations are either computationally expensive (e.g., GAN [69, 45]) or overly simplistic (e.g., eigenbasis [70]). SMAAT uses an alternative approach to find OFM-AEs by perturbing the layer with the lowest ID without the need for manifold approximations. Particularly, SMAAT solves:
Here, $\delta_{l^{*}}$ is the perturbation at the $l^{*}$-th and $\epsilon_{l^{*}}$ is the perturbation budget. $l^{*}$ is the layer with more OFM-AEs composition (lowest ID as assumed in Sec. 4.1). We additionally choose $l^{*}$ to correspond to the layer with the highest index as this would lead to better scalability (shorter Backprop chains). i.e.,
Complexity. Computing the layers IDs and searching the optimal layer $l^{*}$ to perturb are done once per model and per task and incur marginal overhead. For the AT part, when a $P$-step PGD attack is used, this results in $P$ forward-backward passes with length $(n-l^{*}+1)$ instead of $n$. The run-time of every forward/backward pass depends on the layer dimensionality, i.e., $O(d_{l-1}\times d_{l})$ for the $l^{\text{th}}$ layer. Overall, the complexity of one SMAAT forward-backward is $\mathcal{O}((n-l^{*}+1)\max_{l\in[l^{*},n]}(d_{l})^{2}))$. As a result, SMAAT is more efficient than standard AT by a factor of $\mathcal{O}(P\times l^{*}\cdot(\max_{i\in[1,n]}{(d_{i})^{2}}-\max_{j\in[l^{*},n]}{(d_{j})^{2}}))$. In the case of BERT and ROBERT, $l^{*}$ is equal to $n$. The total run-time can be simplified to $\mathcal{O}(P\cdot(d_{n})^{2})$ where $d_{n}$ is the number of classes. Typical enc-LLMs tasks consist of less than five classes [62] which makes the factor of classes small enough to be negligible. Hence, SMAAT enhances the efficiency of the AEs generation process by a factor of $l\cdot O(\max(d_{l}))$. This improvement practically eliminates the cost of the AE generation process.
5 Experiments
We assess SMAAT for its ability to enhance the robustness of various components associated with the use of a language model, aiming to highlight its practical impact. These include: (i) the model itself, particularly in relation to classification tasks; (ii) the retriever model within the Retrieval-Augmented Generation (RAG) setting; (iii) the safeguarding models employed for moderating content produced by generative models. Our robustness evaluation considers four different attacks: (i) word substitution attacks [44, 25, 30], (ii) adaptive attacks [60], (iii) Greedy Coordinate Gradient (GCG) attack [81], and (iv) corpus poisoning attack [78]. We also include a run-time analysis to demonstrate scalability of our method.
5.1 Robustifying Classification Models
Base models. We robustify two base models, BERT-base-cased and RoBERTa-base-cased. Tasks. Considered tasks sentiment clasification (IMDB [36] and YELP [74]) and topic classification (AGNEWS [76]). Attacks. We assess the robustness of models against three input space attacks including PWWS [44] (synonym based), TextFooler [25] (neighbor based), and BERT-Attack [30] (masked language model based). All attacks are conducted using the TextAttack framework [41] and following the settings introduced by Li et al. [32].
Baselines. We compare SMAAT to several baselines including standard (non-adversarial) training, and six baselines from three families of defenses: (1) Input space AT (ASCC [16], DNE [79]), (2) Embedding space AT (FreeLB++ [32]), and (3) Certified defenses (SAFER [71], TMD [40], RSMI [40]). FreeLB++ focuses on scalability and robustness by minimizing the number of PGD steps and applying AT in the initial layer. TMD leverages manifold features by projecting samples back to the manifold in the last layer.
Implementation details. In both base models, $l^{*}=l_{n}$, i.e., we generate AEs in the last layer. Specifically, we perturb the [CLS] embeddings before the classifier layer. We freeze all layers before $l^{*}$ and adversarially train the layers following it for ten epochs with varying attack strength $\epsilon\in[0.1,0.8]$ with learning rate of $0.1$. Further details are presented in Appendix C.
@@ 5.1 Robustifying Classification Models
Table 1: Robustifying sentiment classifiers on AGNEWS, IMDB, and YELP datasets. We report the clean accuracy (CA) and the robust accuracy under attack (AUA) with PWWS (PW), TextFooler (TF), and Bert-Attack (BA), along with the average robust accuracy (AR) across these three attacks. The best performance is bolded.
@@
RSMI 92.7 76.1 63.2 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 92.2 58.7 56.4 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 95.4 45.3 52.3 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA
@@
RSMI 94.3 81.9 74.1 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 93.0 76.2 73.4 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA 96.3 68.9 65.9 NA [^3] As mentioned in the original paper, the current implementation of RSMI takes around 2k times more time than the TextFooler algorithm to generate a single adversarial example with BERT-Attack. Due to this significant time difference, it becomes unfeasible to test RSMI with BERT-Attack. NA
@@
Robustness and generalization.
In Table 1, we report results on robustness and generalization. Note that generalization is high if the adversarially trained models result in higher clean accuracy (CA) than standard training (1st row). On average, SMAAT demonstrates superior robustness over all datasets: it achieves an improvement of 8.6%, 15.7%, and 28.8% over the best score for the BERT model, and 6.0%, 5.8%, and 19.0% for the RoBERTa model on the AGNEWS, IMDB, and YELP datasets, respectively. Note that FreeLB++, which perturbs the first layer, showed the best generalization in four out of six cases, as it produces mostly on-manifold examples (the first layer has a high dimension, as illustrated in Fig. 3) which is inline with the manifold conjecture. SMAAT maintains generalization in five out of six cases and only shows $0.5$ drop in performance in the case of RoBERTa with the IMDB dataset. Besides, SMAAT consistently outperforms TMD, even though it is a manifold based method differently from SMAAT, TMD estimates the manifold and projects the input samples onto it prior to classification. This might lead to inaccuracies in the manifold estimation. SMAAT on the other hand is more robust to such inaccuracies as it only leverage the ID for selecting the layer to perturb. From an attack perspective, RSMI demonstrates superior robustness against PWWS, a synonym-based word substitution attack, in both models. This phenomenon can be attributed to the behavior of the masked inference (MI) component in RSMI, which mimics a synonym-substitution based defense approach without requiring an explicit synonym set, as noted in the original paper [40].
Table 2: Run-time results on IMDB dataset. Mean and standard deviation are computed over ten runs.
Standard ASCC DNE FreeLB++ SAFER RSMI SMAAT

### v2

@@ 3 Notation
Consider a deep neural network classifier $f_{\theta}$ with $n$ layers and parameters $\theta=\{\theta^{(i)}\}_{i=1}^{n}$. We define the transformation spanning layers $l_{i}$ to $l_{j}$ as $f_{\theta^{[i,j]}}=f_{\theta^{(j)}}\circ\dots\circ f_{\theta^{(i)}}$, with $\bar{f}_{\theta^{[i,j]}}$ denoting its standardized version across the dataset $\mathcal{D}$. The input matrix $\mathcal{X}\in\mathbb{R}^{d\times|\mathcal{D}|}$ stacks all samples $\{x_{i}\}_{i=1}^{|\mathcal{D}|}$, while $f_{\theta^{[1,l]}}(\mathcal{X})\in\mathbb{R}^{d_{l}\times|\mathcal{D}|}$ stacks their transformed representations. The data manifold is defined by the high-density region of the sample distribution, distinguishing on-manifold (high-density) from off-manifold (low-density) points. Similarly, the $l^{\text{th}}$ layer manifold is characterized by the density of standardized representations $\{\bar{f}_{\theta^{[1,l]}}(x_{i})\}_{i=1}^{|\mathcal{D}|}$, approximated via the eigenspace of the covariance matrix $\bar{f}_{\theta^{[1,l]}}(\mathcal{X})\bar{f}_{\theta^{[1,l]}}(\mathcal{X})^{T}$. We denote its eigenvectors and eigenvalues as $(U_{l},\Sigma_{l})$, where $U_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ and $\Sigma_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$.
The projection error on the top $k$ eigenvectors is:
$e_{l}^{k}(x)=\bar{f}_{\theta^{[1,l]}}(x)-U_{l}^{k}{U_{l}^{k}}^{T}\bar{f}_{\theta^{[1,l]}}(x),\quad\forall x\in\mathcal{X}.$ (2)
The eigen-based manifold dimension $k_{l}$ is the minimum $k$ required for a reconstruction error $\gamma$:
$k_{l}=\min_{k}\left\{k\in[1,d_{l}]\;\Big{|}\;\sum_{x\in\mathcal{D}}\|e_{l}^{k}(x)\|_{2}\leq\gamma\right\}.$ (3)
We classify adversarial examples (AEs) based on projection error: - A ($\gamma,l$)-off-manifold AE (OFM-AE) has $e_{l}^{k}>\gamma$. - A ($\gamma,l$)-on-manifold AE (ONM-AE) has $e_{l}^{k}\leq\gamma$. Finally, we denote the Intrinsic Dimension (ID) of layer $l$, estimated via TwoNN, as $I_{l}$, and its normalized ID as $I_{l}^{d}=I_{l}/d_{l}$.
4 Exploring Layerwise ID and Its Effects
Our study is motivated by the not-well understood difference in the generalization and robustness trends between adversarially trained vision models and LLMs. Specifically, (1) robustness improvements are consistently more pronounced in vision models (Madry et al., 2017; Altinisik et al., 2023b), (2) generalization either remains stable or improves in enc-LLMs (Tsipras et al., 2019; Zhu et al., 2020), whereas it deteriorates in vision models, and (3) dec-LLMs achieve the optimal trade-off when AT is applied across multiple network layers (Casper et al., 2024; Sheshadri et al., 2024).
We demonstrate that the varying trends observed across different network architectures are closely linked to the layer-wise distribution of OFM-AEs. Notably, we find that in the first layer, where conventional AT is applied, adversarial examples for enc-LLMs tend to be more on-manifold, while vision models and dec-LLMs show a higher proportion of off-manifold examples (see Fig. 3, left). This explains the generalization and robustness trends as, by the manifold conjecture, higher proportions of ONM and OFM AEs lead, respectively, to better generalization and robustness (Ma et al., 2018; Stutz et al., 2019a; Alemany & Pissinou, 2020; Li et al., 2021a).
(a) LLaMA-2
(b) BERT
(c) VGG
Figure 5: Layer-wise effects of adversarial training on robustness and generalization across model types. Each subplot shows the impact of applying AT at different layers of (a) LLaMA-2 (dec-LLM), (b) BERT (enc-LLM), and (c) VGG (vision model). Marker colors transition from light blue (lower layers) to dark blue (higher layers). Observed trends align with changes in intrinsic dimensionality and the distribution of on- vs. off-manifold adversarial examples, as shown in Figure 4.
ID & ONM/OFM Relationship We hypothesize that lower intrinsic dimensionality ($I_{l}^{d}$) leads to a higher proportion of OFM-AEs, whereas higher $I_{l}^{d}$ yields more ONM-AEs. Formally, let $\mathcal{M}\subset\mathbb{R}^{n}$ be a smooth, compact, low-dimensional manifold with intrinsic dimension $I_{l}^{d}\ll n$, embedded in $\mathbb{R}^{n}$. Let $f:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be a classifier trained on data sampled from $\mathcal{M}$. Since the classifier learns $p(y\mid x)$ without access to the true generative distribution $p(x)$, it lacks explicit knowledge of the underlying data manifold $\mathcal{M}$. Consequently, the loss gradient $\nabla_{\delta}\mathcal{L}$ often contains components that are orthogonal to the tangent space of $\mathcal{M}$. As a result, the perturbation $\delta$ is unlikely to remain entirely within the manifold and typically includes a non-zero component in the orthogonal direction. Moreover, when $I_{l}^{d}$ is small relative to $n$, the proportion of $\delta$ that lies off the manifold increases. Thus, adversarial examples generated via gradient-based methods are more likely to fall outside the data manifold, i.e., to be off-manifold adversarial examples.
We base this conclusion on an empirical investigation into the relationship between the ID and the proportion of ONM/OFM AEs across different layers of various deep neural network models, including two enc-LLMs (BERT and RoBERTa), three vision models (ViT (Dosovitskiy et al., 2020), DeiT (Touvron et al., 2021), VGG-13 (Simonyan & Zisserman, 2014)), and three dec-LLMs (LLaMA-2-7b-chat (Touvron et al., 2023), Vicuna (Zheng et al., 2024), Mistral-7b-instruct (Jiang et al., 2023)) on several datasets (refer to Appendix B for details). Results are reported in Figure 4. The ID ($I_{l}^{d}$) in row 1, is computed using twoNN estimation(Facco et al., 2017) and follows the OFM-AEs trend (row 2). To assess the proportion of OFM/ONM AEs across different layers, we use the reconstruction error $e^{k}_{l}$ at layer $l$ (row 2) as a metric, i.e., higher projection error corresponds to higher proportion of OFM-AEs.
ID & Robustness/Generalization Relationship
We extend our evaluation of the relationship between intrinsic dimensionality (ID), robustness, and generalization by applying adversarial training (AT) to different layers of dec-LLMs (LLaMA-2-7B-chat), enc-LLMs (BERT), and vision models (VGG). For the dec-LLM, we train LLaMA-2-7B-chat on the LAT dataset (Sheshadri et al., 2024) using various hyperparameter configurations (see Appendix C for details), applying LAT at every even-numbered layer. To evaluate generalization, we use MT-Bench (Zheng et al., 2023), while robustness is assessed via failure rates against the GCG attack , which generates adversarial suffixes that bypass language model safety alignment. Additional results using the PAIR (Chao et al., 2023) attack are provided in Appendix C. As shown in Fig. 5(a), applying LAT to lower layers (light blue markers) enhances robustness but reduces generalization (bottom-right), while applying it to upper layers (dark blue markers) improves generalization at the cost of robustness (top-left). This trend aligns with the intrinsic dimension (ID), which decreases in higher layers, as well as with the distribution of ONM/OFM adversarial examples. These results also explain the superiority of the LAT approach of Sheshadri et al., 2024 which applies AT to four evenly spaced layers across the network.
We conducted a similar experiment on the BERT model using the YELP dataset (see Sec. 6.1 for details). Our results empirically show that, following the decreasing ID trend, applying AT at higher layers leads to increased robustness as the layer index $i$ increases. Figure 5(b) presents the results of AT on BERT for the YELP dataset. As expected, robustness (measured under the TextFooler attack) improves with higher layer indices while generalization remains unaffected.
@@ 4 Exploring Layerwise ID and Its Effects
Beyond providing an explanation for the difference in robustness and generalization trends in vision models, enc- and dec-LLMs, results of Fig. 4 shed light on an interesting trend for the ID across layers. While the ID is increasing in vision and dec-LLMs, it is monotonically decreasing in enc-LLMs (Bert and RoBERTa). This motivated the design of a new algorithm that controls the gains in robustness and generalization via controlling the proportions of ONM- and OFM-AEs. This can be achieved by generating AEs in the intermediate layers of the deepnet based on their IDs. To this end, we propose SMAAT (See Fig. 3, right), an AT algorithm that aims at achieving high robustness by perturbing the layer $l^{*}$ with the lowest ID to generate a high proportion of OFM-AEs. A side effect of perturbing an intermediate layer as opposed to the input one is significantly improving scalability by a factor $\mathcal{O}\Big{(}P(n-l^{*})\Big{(}\max\!\Big{(}\{d_{l_{i}}|l_{i}\in[1,l^{*}-1]\}\Big{)}-d_{l^{*}}\Big{)}\Big{)}$ as it results in shorter $P$-step PGD chains.
@@
Table 1: Robustifying sentiment classifiers on AGNEWS, IMDB, and YELP datasets. We report the clean accuracy (CA) and the robust accuracy under attack (AUA) with PWWS (PW), TextFooler (TF), and Bert-Attack (BA), along with the average robust accuracy (AR) across these three attacks. The best performance is bolded. Across all data sets, SMAAT achieves high robustness (AR) while maintaining high generalization (CA).
@@
RSMI 92.7 76.1 63.2 NA [^2] RSMI takes about 2k times longer than TextFooler to generate a AE with BERT-Attack, making it unfeasible to test. NA 92.2 58.7 56.4 NA [^2] RSMI takes about 2k times longer than TextFooler to generate a AE with BERT-Attack, making it unfeasible to test. NA 95.4 45.3 52.3 NA [^2] RSMI takes about 2k times longer than TextFooler to generate a AE with BERT-Attack, making it unfeasible to test. NA
@@
RSMI 94.3 81.9 74.1 NA [^2] RSMI takes about 2k times longer than TextFooler to generate a AE with BERT-Attack, making it unfeasible to test. NA 93.0 76.2 73.4 NA [^2] RSMI takes about 2k times longer than TextFooler to generate a AE with BERT-Attack, making it unfeasible to test. NA 96.3 68.9 65.9 NA [^2] RSMI takes about 2k times longer than TextFooler to generate a AE with BERT-Attack, making it unfeasible to test. NA
@@
5 SMAAT: Scalable Manifold Aware Adversarial Training
We propose SMAAT an AT algorithm that aims at improving the robustness and scalability of standard AT by generating AEs in the layer leading to the highest proportion of OFM instead of from the input layer as classically. Specifically, given a pretrained model $f_{\theta}$, SMAAT intentionally generates a higher proportion of OFM-AEs to enhance robustness. Formally, SMAAT solves the augmented AT objective:
[…]
While the manifold conjecture refers to input space AEs, we relax it to encompass intermediate layers as well, i.e., we chose the perturbation $\delta$ that results in an OFM-AE at any layer across the deepnet. Intuitively, AEs that are off- or on- the transformed data manifold in any layer also affect robustness. The relaxed objective becomes:
[…]
Note that solving the above objective using the method of Lagrangian Multipliers is possible, but it would require an approximation of the manifold to characterize OFM-AEs. Such approximations are either computationally expensive (e.g., GAN (Xiao et al., 2018)) or overly simplistic (e.g., eigenbasis (Xiao et al., 2022)). SMAAT uses an alternative approach to find OFM-AEs by perturbing the layer with the lowest ID without the need for manifold approximations. Particularly, SMAAT applies AT at $l^{*}$-th layer where the layer with more OFM-AEs composition (lowest ID):
We additionally choose $l^{*}$ to correspond to the layer with the highest index as this would lead to better scalability (shorter PGD chains). as well as to potentially the highest proportion of off-manifold AEs at any intermediate layer, i.e.,
Table 2: Run-time results on IMDB dataset. Mean and standard deviation are computed over ten runs. Compared to FreeLB++, training time of SMAAT is lower by a factor of nearly three.
Standard ASCC FreeLB++ SAFER RSMI SMAAT